OpenStudy (anonymous):

–(–33 + 18) = 33 + (–18) What property is shown by the equation?

OpenStudy (anonymous):

2. –10 + (–15) = –15 + (–10) and these too 3. 4 + (29 + 16) = 4 + (16 + 29) 4. –12 + (12 + 27) = (–12 + 12) + 27